OpenStudy (flygirlmandi):

2 1/7 times 3/4 ? a-3 3/28 b-2 1/7 c-1 17/28 d-2 3/28 thanks :)

OpenStudy (yoloshroom):

2(1/7) x 3/4?

MsBrains (ms-brains):

C @flygirlmandi Medal me :)

OpenStudy (flygirlmandi):

are you sure

OpenStudy (yoloshroom):

Brains that's not how we do it. Explain it out :)

MsBrains (ms-brains):

yes. I am.

MsBrains (ms-brains):

I positive!

OpenStudy (tkhunny):

Please convert the first to an improper fractions.

OpenStudy (flygirlmandi):

15/7 @tkhunny

OpenStudy (tkhunny):

Yes. Multiply away! \(\dfrac{15}{7}\cdot\dfrac{3}{4}\)